The quality control manager at a computer manufacturing company believes that the mean life of a computer is 117 months with a standard deviation of eight months if he is correct what is the probability that the mean of a sample of 52 computers would be less than 115. 46 months round your answer to four decimal places
Answer #1

ANSWER:

Given that,

The quality control manager at a computer manufacturing company believes that the mean life of a computer is 117 months with a standard deviation of eight months if he is correct what is the probability that the mean of a sample of 52 computers would be less than 115. 46 months round your answer to four decimal places

Mean = \mu = 117

Standard deviation = s = 8

Sample size = n= 52

P(\bar{x} < 115.46) = P((\bar{x}-\mu)/(s/sqrt(n)) < (115.46-117)/(8/sqrt(52)))

P(\bar{x} < 115.46) = P(Z < (-1.54)/(8/sqrt(52)))

P(\bar{x} < 115.46) = P(Z < -1.39)

P(\bar{x} < 115.46) = 0.08226 (From z score table)

P(\bar{x} < 115.46) = 0.0823 (Rounded to four decimal places)

STANDARD NORMAL DISTRIBUTION: Tod values Regrosat AREA wo he LEFT Stiho Zsore, na Z -3.8 -3.7 -3.4 -3.3 -3.0 -2.9 STANDARD NO

-----------------------------------------------The End ------------------------------------------------------------------------

If you have any doubt please ask me friend and if you are satisfied with my solution please give me thumb up.